A cold water faucet can fill a sink within 12 minutes and a hot water faucet can fill it in 15 minutes. The drain can empty the sink in 25 minutes. If both faucets are on and the drain is open, how long will it take to fill the sink?
Found 2 solutions by amoresroy, piyushchy:
Answer by amoresroy(333)

Let t = the number of minutes it will take to fill the sink
t(1/12 + 1/15 - 1/25) = 1
t (25/300 + 20/300 - 12/300) = 1
33t/300 = 1
t = 300/33
t = 9.09 mins

Answer by piyushchy(1)

Let t = the number of minutes it will take to fill the sink
t(1/12 + 1/15 - 1/25) = 1
t (25/300 + 20/300 - 12/300) = 1
33t/300 = 1
t = 300/33
t = 9.09 mins